find the sum of all the natural numbers between 1 and 200 which are neither divisible by 2 nor5

Grade:11

2 Answers

gajanana hegde
10 years ago

sum of odd terms from 1to 199

number of terms is =100 (tn=a +(n-1)d)a=1;d=2;tn=199

Sn(odd)=n/2[a+l]=50*200=10000

now deduct the term which are multiples of 5

a=5

d=10

tn=>195=5+(n-1)10

n=10

Sn=n/2[a+l]

5*200=1000

.:.total=10000-1000=9000

srinivas aytha
10 years ago

1+2+............+200=n/2(a+l)=200/2(1+200)=100(201)=20100-10100(div by 2)=10000, 10000-(5+15+......195)=10000-20/2(5+195)=10000-2000=8000